[Brahms-l] Shift report 20050126 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Wed Jan 26 2005 - 23:53:00 EST
There were problems with the beam at the beginning after the initial store was dumped.  Several aborted Fills and ramps.  Had beam since about 9:00.  Following runs achieved
run       tr2         tr3
12514     94k        34k
12522     end run died, so this information unavailable
12523    186k        67k
12524    100k        37k     This run still in progress

Noticed that the hall probe value of D4 is near 0.  It is probably not a problem with analysis, but should probably be checked.

Hiro came and fixed several channels of Si that had been absent.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed Jan 26 23:53:29 2005

This archive was generated by hypermail 2.1.8 : Wed Jan 26 2005 - 23:53:53 EST